OpenStudy (samirahdanyel):

The volume of an object is equal to the ratio of its mass to density, V = . The mass of a spherical grape is 8.4 grams and its density is 2 grams per cubic centimeter. What is the radius of the grape? Round to the nearest tenth of a centimeter. 1.0 cm 1.5 cm 1.9 cm 2.1 cm

OpenStudy (samirahdanyel):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`The volume of an object is equal to the ratio of its mass to density, V = . The mass of a spherical grape is 8.4 grams and its density is 2 grams per cubic centimeter.` your equation got cut off. It just says `V = ` and nothing after the equal sign

OpenStudy (samirahdanyel):

Okay hold on

jimthompson5910 (jim_thompson5910):

ok so first we need the volume V m = 8.4 d = 2 V = ??? use the formula given

OpenStudy (samirahdanyel):

its 4.2

jimthompson5910 (jim_thompson5910):

yes V = 4.2

jimthompson5910 (jim_thompson5910):

This is the volume of a sphere. We now turn to the formula for the volume of a sphere \[\Large V = \frac{4}{3}\pi*r^3\] something you should have in your notes, on a notecard, on a reference sheet, or have it memorized

jimthompson5910 (jim_thompson5910):

plug in V = 4.2 and then solve for r

OpenStudy (samirahdanyel):

yeah i know

OpenStudy (samirahdanyel):

i got 98.784 pi

jimthompson5910 (jim_thompson5910):

\[\Large V = \frac{4}{3}\pi*r^3\] \[\Large 4.2 = \frac{4}{3}\pi*r^3\] \[\Large 3*4.2 = 3*\frac{4}{3}\pi*r^3\] \[\Large 12.6 = 4\pi*r^3\] \[\Large \frac{12.6}{4\pi} = \frac{4\pi*r^3}{4\pi}\] \[\Large \frac{12.6}{4*3.14} = r^3\] I'll let you finish up

OpenStudy (samirahdanyel):

What do I put in the calcuator

jimthompson5910 (jim_thompson5910):

you need to evaluate the left side first, so evaluate `12.6/(4*3.14)`

OpenStudy (samirahdanyel):

i got 3.15/pi

jimthompson5910 (jim_thompson5910):

now use a calculator to evaluate 3.15/pi

OpenStudy (samirahdanyel):

thank you

jimthompson5910 (jim_thompson5910):

you then have to take the cube root of both sides but it sounds like you found the answer already